In a WAITS, one subtest the raw score has a mean of 55 and variance of 25. What is the probability of obtaining a raw score between the mean and 46? AND between 50 and 58?
1
Expert's answer
2020-05-31T17:39:52-0400

"Given \; that, μ=55, σ^{2}=25, σ=5,then,\\
 a) P(46<x<55) = P(\frac{46-55}{5}< Z <\frac{55-55}{5})\\
=P(-1.8<Z<0)\\= P(0<Z<1.8)=0.4641\\
 b) P(50<x<58) = P(\frac{50-55}{5}< Z <\frac{58-55}{5})\\
=P(-1<Z<0.6)\\
= P(0<Z<1)+ P(0<Z<0.6)\\
=0.3413+0.2257=0.567"